# VideoChat-Flash: Hierarchical Compression for Long-Context Video Modeling

### v2

@@ 5 Conclusions
In this paper, we address the challenge of long-context video modeling in MLLMs by introducing the Hierarchical video token Compression (HiCo). HiCo effectively reduces the compute through a hierarchical strategy capitalizing redundancies in visual details and semantics, preserving vast amounts of visual essentials in a compact form. Based on HiCo, we design a practical context modeling system VideoChat-Flash. With its multi-stage short-to-long learning scheme, rich dataset of real-world long videos, and robust distributed computing framework, VideoChat-Flash ensures that MLLMs can be trained and deployed at scale for handling hour-scale videos with millions of tokens, achieving leading performance in long and short video benchmarks.
@@
6 Implementation Details and More Analysiss
6.1 Clip Compressor
@@ 6.2 Visual Dropout in LLM
7 Training Details

### v4

@@ 5 Conclusions
In this paper, we address the challenge of long-context video modeling in MLLMs from the model architecture, training data, training strategy and evaluation benchmark. We design an efficient architecture for video MLLMs by introducing a hierarchical long video context compression method, which achieves an extreme compression ratio with nearly no performance loss. Regarding data and training, we propose a new long video training corpus and short-to-long learning strategy, which effectively enhances the model’s understanding ability for videos of various lengths. Additionally, we developed a new and more challenging long video context evaluation benchmark. Our model demonstrated outstanding performance on various video understanding benchmarks, which validates the effectiveness of our proposed methods.
@@
6 More Results & Discussions
@@ 6.1 Visual Dropout in LLM
Video encoder MVBench PerceptionTest LongVideoBench MLVU VideoMME (w/o sub.) LVBench Avg Val Val M-Avg Overall Avg Avg. Duration 16s 23s 473s 651s 1010s 4101s UMT-L 73.2 75.6 64.2 74.5 64.0 48.4 InternVideo2-1B 74.(+1.1) 76.3(+0.7) 64.5(+0.3) 73.4(-1.1) 65.2(+1.2) 48.7(+0.3)
Table 7: Results with different video encoder.
6.2 Results with InternVideo2
As shown in Tab. 7, in addition to UMT [28], we also attempted to use the more powerful InternVideo2-1B [56] as the video encoder. As shown in Table 1, we found that a stronger video encoder can lead to better compressed representations.
6.3 Results on Image Understanding Benchmarks
Our model is specifically designed for video understanding. However, according to the newly-evaluated results of image benchmarks, our model can still outperform the strong image-based MLLM, LLaVA-NeXT [70], with significantly lower computational cost: MMMU (45.2 vs. 35.3), MME (1843.4 vs 1603.7).
7 Implementation Details
7.1 Video-Language Connectors
Figure 10: Comparison of different connectors.
@@ 8.4 HD-VILA
9 Qualitative Results
We perform qualitative comparisons of our model with the proprietary model Gemini-1.5 Pro [44] [^1] We use the newest Gemini-1.5 Pro-002 for evaluation. and the open-source LongVU [47] and VideoLLaMA2 [14] across three tasks: fine-grained understanding of short videos ( Figs. 11 and 12) and long video understanding ( Figs. 13 and 14).
Figure 11: Fine-grained Understanding of Short Videos: Fast Motion. By adopting a dense sampling strategy for short videos, our model effectively captures fast motion within the video, enabling it to accurately determine the final position of the object under the cup.
Figure 12: Fine-grained Understanding of Short Videos: Action Count. Our model can accurately capture actions in short videos while also recording their occurrence frequency.
Figure 13: Long video understanding: Anomaly Detection In Surveillance Videos. Our model can detect anomalous behaviors in surveillance videos and provide corresponding inferences.
Figure 14: Long video understanding: Moive Understanding. Our model can understand the plot of a movie and retain detailed visuals.